So I have to tell that my english isn't to good , so i will be able to explain exacly what i need if that will be needed
1) how big power do you need to use to make 1 kg body fly up to 50 meters ?
1
Expert's answer
2011-10-20T12:13:31-0400
First let's find initial velosity of the body required to reach 50m
energy conservation law:
mv^2/2=mgh where m -mass of the body, v - initial velosity, h=50m
hence v=sqrt(2gh)=31 m/s
When making body to move force will do work A=mv^2/2/ where v -initial velosity
So A=1*31^2/2=490,5 joules -& work done to make body fly up to 50m